OpenStudy (impracticaljoker):

4) How many solutions does each equation have a) 4y – 4y – 12 = 14 – 2 b) –3y + 3y + 4 = 4 c) 3x + 6 = –1 – 3 + 4x

OpenStudy (impracticaljoker):

@rebeccaxhawaii

OpenStudy (impracticaljoker):

@FortyTheRapper

OpenStudy (impracticaljoker):

@rebeccaxhawaii

OpenStudy (impracticaljoker):

Dont you dare leave me ;c

OpenStudy (impracticaljoker):

@rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

im sorry ;c

OpenStudy (impracticaljoker):

Heeelp

OpenStudy (impracticaljoker):

Go to that link?

OpenStudy (fortytherapper):

A.) and B.) are going to be the same. What do you notice about the Y values?

OpenStudy (impracticaljoker):

What do you mean? I see there are 2 in both equations with the same numbers.

OpenStudy (fortytherapper):

That's a good start Are the both positive, negative, or one of each?

OpenStudy (impracticaljoker):

One of each

OpenStudy (fortytherapper):

Right, so for the first one \[4y-4y=?\]

OpenStudy (impracticaljoker):

Wait, I dont understand what you're asking for

OpenStudy (fortytherapper):

Just a subtraction problem If you had "4y" of something and I took away "4y" of something from you, how much would you have

OpenStudy (anonymous):

0

OpenStudy (impracticaljoker):

just the y or would that be gone too?

OpenStudy (fortytherapper):

Right, it would be gone too because it works the same way as just numbers. If it helps, you could forget about the y 4-4 = 0 Then add the y back on 0y, which is 0 basically

OpenStudy (impracticaljoker):

Ok

OpenStudy (impracticaljoker):

Next step?

OpenStudy (fortytherapper):

Which means the y is gone, leaving us with \[-12=14-2\] (Sorry, Im having some script problems so I'm lagging)

OpenStudy (fortytherapper):

If that statement true or false

OpenStudy (impracticaljoker):

So no solution1

OpenStudy (fortytherapper):

Correct

OpenStudy (impracticaljoker):

And since the 2nd equation is the same it will be no solution for that one too!

OpenStudy (fortytherapper):

Not exactly the same, but it is the same to an extent of the y's -3y + 3y is saying -3+3=? That is 0, so that leaves us with 4=4 That's true or false?

OpenStudy (impracticaljoker):

Not understanding what you're asking again.

OpenStudy (fortytherapper):

I am back from work Part B: \[–3y + 3y + 4 = 4 \] When we subtracted 3y-3y, we got 0, giving us \[4 = 4\] This means the line of y = 4 is on top of the line of y = 4, so for whatever value X we look at, the y value for both lines will always be the same... which would be 4. That means we have an infinite number of solutions.